# Site Basis Excitation Ansatz for Matrix Product States

### v1

@@ Abstract
The spectrum of elementary excitations of a quantum system is one of its most important properties. For a one-dimensional system, conventional density matrix renormalization group (DMRG)[1, 2] methods allow one to obtain a limited number of excited states to be obtained, but obtaining a full dispersion relation is difficult. Dynamical algorithms, such as dynamical DMRG[3], which works in frequency space, or those that use real-time evolution[4, 5, 6, 7, 8, 9], offer an alternative[10], but a dramatic increase in efficiency arose with the introduction of the matrix product state (MPS) version of the single-mode approximation, called the excitation ansatz (EA)[11, 12]. In EA, the excitation is a plane-wave superposition of MPS, where in each MPS the tensor on one site is different from that of the ground state and creates a local excitation near that site. Remarkably, the accuracy of this simple ansatz for simple gapped systems often rivals that of DMRG itself.
@@
where $\mathbf{S}_{j}$ are the spin-1 operators at site $j$ and $J=1$. This model is gapped[14, 15], and the lowest excitations are triplet magnons[16, 17, 18]. Two of its features, being an integer spin chain and one with a finite gap and correlation length ($\xi\approx 6$), are important for the algorithm used. The finite correlation length means that finite system DMRG can reach the bulk in the center of a finite system. The $\ldots AAA\ldots$ single-site infinite MPS form is appropriate for integer spin chains. (Here and below, we will frequently use this simplified notation for uniform MPS, which indicates the unit cell.) Note that we will make use of the finite (and modest) correlation length to utilize certain algorithmic simplifications—shortcuts, which are not central to SBEA, but which may be very convenient in avoiding some technical complications particularly for the non-expert.
@@
We start by performing finite system DMRG on a long open chain ($N\gg\xi$, e.g. $N=150$), where a matrix product operator (MPO) form for the Hamiltonian is convenient, particularly since MPOs can be generated automatically[20]. For $N\gg\xi$ the center is representative of the infinite system in its local properties, but the central tensors have random gauges, making them unsuitable for making an infinite uniform MPS. Specifically, we can say the states of the links on each side of each tensor are different, whereas for the $\ldots AAA\ldots$ form they should be the same.
To make the links the same, we insert a new site in the middle, with a new tensor. (Inserting sites in the middle goes all the way back to the original infinite-system DMRG[1, 2].) To accommodate the new $N+1$ sites, we simply reconstruct the Hamiltonian MPO for $N+1$ sites. Picking a central link on the $N$-site system, we put the orthogonality center there, and obtain a $\tilde{\lambda}$ diagonal matrix of Schmidt coefficients for that link. (For example, to put it between tensors $CD$, we contract these two tensors and then split them apart with a singular value decomposition, obtaining $C^{\prime}\tilde{\lambda}D^{\prime}$.) The half-MPS on either side of the link form orthogonal bases for $\tilde{\lambda}$. We add a new site at that link, replacing $\tilde{\lambda}$ with a new three leg tensor $\tilde{D}$, which we initialize randomly. We then perform a Lanczos diagonalization to optimize the energy, to high accuracy, equivalent to a single step of single site finite-size DMRG, but without a good initial state. Since this is done only once, the extra cost of the poor initialization and high accuracy is negligible. As in the original infinite system DMRG, the overall state obtained (assuming a single site unit cell is appropriate) is just as accurate for $N+1$ sites as the state it started from was on $N$ sites. And, importantly, $\tilde{D}$ has identical left and right indices, so it can immediately make an infinite MPS.
@@
The site basis excitation ansatz.— In the EA[24, 11], (with unit cell size $p=1$) we consider infinite MPS where one $A$ tensor has been replaced by another, $B$. If this is done at site $j$, we denote this as $|B\rangle_{j}$. The EA for an excitation with momentum $k$ is
@@
where $B$ is optimized for the minimum energy separately at each $k$, so that we have $B(k)$, and we obtain an excitation energy $E(k)$. Although this is a superposition of MPS, one for each possible position of $B$, it was observed by Osborne and McCulloch [23] that this state can also be represented exactly as a single MPS with doubled bond dimension.
@@
Instead, we make an additional ansatz for the $B_{\alpha}$}: we obtain them from a single site Lanczos calculation of a single $B$ on one site, similar to that used above in the generation of the infinite ground state, but we use it to obtain lowest $N_{\alpha}$ eigenstates of the effective Hamiltonian for the single site, where $N_{\alpha}$ is set to go up to an energy cutoff. Why could this be a good ansatz? The degrees of freedom involved are the states of the site and the Schmidt vectors on either side of the site, extending into the bulk. The highest probability Schmidt vectors are quite local to the site; lower probability vectors extend farther and farther into the bulk. If we imagine the Schmidt vectors for a finite bond dimension of the ground state extending a certain distance into the bulk, then as we find more eigenstates we are approximating a complete set of states in the region near the site, but with an energy cutoff. Then the EA superposes that group of states over all sites, so it is natural to expect a nearly complete basis overall up to the energy cutoff. However, the states involved all have a single $B$, so we expect the ansatz could cover the single excitation space but have the same limitations as EA, e.g. not covering two-particle excitations. In addition, there will be substantial overlap of these states from nearby sites, so in constructing the full basis over all the sites it is essential to treat the non-orthogonality properly.
@@
Figure 2: Spin profile of single site excitations $\langle S_{z}(x)\rangle$ versus site $x$. (a) Excitation profiles for the lowest energy single-site state, where a single tensor is different from the ground state. The numbers beside each curve indicate the cutoff $\lambda^{2}$ for the pretruncation step; these cutoffs result in $\chi_{\text{tr}}=8,24,46$, and 9090 for the curves top to bottom. (b) Comparison of excitation profile for states with and without restriction to the left null space, both with $\lambda^{2}=10^{-10}$.
@@
Because the ground state is gapped, connected correlations decay as $e^{-|j-j^{\prime}|/\xi}$ and both kernels are therefore exponentially banded. In practice we take $L_{c}\leq j-j^{\prime}\leq L_{c}$ with $L_{c}\simeq 150$; beyond that threshold every entry is below about $10^{-12}$ and can be safely discarded. The calculation of the $H$ array of size $N_{\alpha}^{2}L_{c}$ is the dominant effort in SBEA. If the Hamiltonian MPO is in infinite uniform $\ldots HHH\ldots$ form, then the contractions are particularly simple, since the edge tensors don’t depend on the site. In that case for each $B_{\alpha}$ we start from the left, and progressive contract towards the right. For each site, there is a final contraction of all the $B_{\alpha^{\prime}}$. Then the calculation time is $O[(N_{\alpha}^{2}+N_{\alpha}K)L_{c}\chi^{3}]$, where $K$ is the MPO bond dimension. This calculation time is identical to that of a finite-system DMRG calculation on a system of length $L_{c}$, except the parenthetical factor would be replaced by an effective number of sweeps times the number of Lanczos steps at each step.

### v2

# arXiv:2509.06241v2[cond-mat.str-el] 02 Mar 2026

@@ Abstract
The spectrum of elementary excitations of a quantum system is of fundamental importance both for theoretical understanding and for direct connection with experiments. Spectral functions are generally more difficult to obtain compared to static properties, and a number of numerical approaches have been utilized in the past, such as small-system exact diagonalization combined with Lanczos/continued-fraction evaluation of dynamical correlators[5, 6], or cluster perturbation theory, which extends exact diagonalization with perturbation theory[20, 21]. For a one-dimensional system, conventional density matrix renormalization group (DMRG)[29, 30] methods allow one to obtain a limited number of excited states to be obtained, but obtaining a full dispersion relation is difficult. Dynamical algorithms, such as dynamical DMRG[12], which works in frequency space, or those that use real-time evolution[25, 3, 27, 7, 19, 18], offer an alternative[26]. There is current great interest in applying such methods to more challenging two-dimensional systems, which may be at the limits of current computational capability. A dramatic increase in efficiency arose with the introduction of the matrix product state (MPS) version of the single-mode approximation, called the excitation ansatz (EA)[9, 23]. In EA, the excitation is a plane-wave superposition of MPS, where in each MPS the tensor on one site is different from that of the ground state and creates a local excitation near that site. Remarkably, the accuracy of this simple ansatz for simple gapped systems often rivals that of DMRG itself.
@@
where $\mathbf{S}_{j}$ are the spin-1 operators at site $j$ and $J=1$. This model is gapped[11, 10], and the lowest excitations are triplet magnons[1, 2, 28]. This well-understood model serves as a benchmark and development platform, rather than a system where we obtain new insight. We hope that our developments will serve to facilitate studies of more difficult systems, particularly in two dimensions. Two of the model’s features, being an integer spin chain and one with a finite gap and correlation length ($\xi\approx 6$), are important for the algorithm used. The finite correlation length means that finite system DMRG can reach the bulk in the center of a finite system. The $\ldots AAA\ldots$ single-site infinite MPS form is appropriate for integer spin chains. (Here and below, we will frequently use this simplified notation for MPS, where here the notation indicates that all of the tensors are identical in a uniform MPS. It is to be understood that the tensors $A$ are matrices in their connections to their neighbors, but that they also have a third index, which denotes the physical states of a site.) Note that we will make use of the finite (and modest) correlation length to utilize certain algorithmic simplifications—shortcuts, which are not central to SBEA, but which may be very convenient in avoiding some technical complications particularly for the non-expert.
@@
We start by performing finite system DMRG on a long open chain ($N\gg\xi$, e.g. $N=150$), where a matrix product operator (MPO) form for the Hamiltonian is convenient, particularly since MPOs can be generated automatically[4]. For $N\gg\xi$ the center is representative of the infinite system in its local properties, but the central tensors have random gauges. The random gauges stem from the fact that inserting on any link $MM^{-1}$, with $M$ an ordinary nonsingular matrix, leaves the wavefunction unchanged. Left or right orthogonality of each tensor, which DMRG requires away from the orthogonality center, does not resolve this gauge ambiguity, and the $\ldots ABCD\ldots$ form interferes with forming the desired $\ldots AAA\ldots$ infinite uniform MPS. Specifically, we can say the states of the links on each side of each tensor are different, whereas for the $\ldots AAA\ldots$ form they should be the same.
To make the links the same, we insert a new site in the middle, with a new tensor. (Inserting sites in the middle goes all the way back to the original infinite-system DMRG[29, 30].) To accommodate the new site, we simply reconstruct the Hamiltonian MPO for $N+1$ sites. Picking a central link on the $N$-site system, we put the orthogonality center there, and obtain a $\tilde{\lambda}$ diagonal matrix of Schmidt coefficients for that link. (For example, to put it between tensors $CD$, we contract these two tensors and then split them apart with a singular value decomposition, obtaining $C^{\prime}\tilde{\lambda}D^{\prime}$.) The half-MPS on either side of the link form orthogonal bases for $\tilde{\lambda}$. We add a new site at that link, replacing $\tilde{\lambda}$ with a new three leg tensor $\tilde{D}$, which we initialize randomly. We then perform a Lanczos diagonalization to optimize the energy, to high accuracy, equivalent to a single step of single site finite-size DMRG, but without a good initial state. Since this is done only once, the extra cost of the poor initialization and high accuracy is negligible. As in the original infinite system DMRG, the overall state obtained (assuming a single site unit cell is appropriate) is just as accurate for $N+1$ sites as the state it started from was on $N$ sites. And, importantly, $\tilde{D}$ has identical left and right indices, so it can immediately make an infinite MPS.
@@
The site basis excitation ansatz.— In the EA[8, 9], (with unit cell size $p=1$) we consider an infinite MPS where one ground-state $A$ tensor has been replaced by another tensor, $B$, i.e. a $\ldots AAABAAA\ldots$ form. If the $B$ is at site $j$, we denote this as $|B\rangle_{j}$. The EA for an excitation with momentum $k$ is
@@
where the tensor $B$ is the same no matter which site it is on. This $B$ is determined by optimizing the excitation for the minimum energy, separately for each $k$, so that we have $B(k)$, and we obtain an excitation energy $E(k)$. Although there is only one $B$ to determine for each $k$, the optimization via an effective Hamiltonian is more complicated than in DMRG where each site tensor is independent. Although this is a superposition of MPS, one for each possible position of $B$, it was observed by Osborne and McCulloch [17] that this state can also be represented exactly as a single MPS with doubled bond dimension.
@@
Instead, we make an additional ansatz for the $B_{\alpha}$}: we obtain them from a single site Lanczos calculation of a single $B$ on one site, similar to that used above in the generation of the infinite ground state, but we use it to obtain the lowest $N_{\alpha}$ eigenstates of the effective Hamiltonian for the single site, where $N_{\alpha}$ is set to go up to an energy cutoff. Why could this be a good ansatz? The degrees of freedom involved are the states of the site and the Schmidt vectors on either side of the site, extending into the bulk. The highest probability Schmidt vectors are quite local to the site; lower probability vectors extend farther and farther into the bulk. If we imagine the Schmidt vectors for a finite bond dimension of the ground state extending a certain distance into the bulk, then as we find more eigenstates we are approximating a complete set of states in the region near the site, but with an energy cutoff. Then the EA superposes that group of states over all sites, so it is natural to expect a nearly complete basis overall up to the energy cutoff. However, the states involved all have a single $B$, so we expect the ansatz could cover the single excitation space but have the same limitations as EA, e.g. not covering two-particle excitations. In addition, there will be substantial overlap of these states from nearby sites, so in constructing the full basis over all the sites it is essential to treat the non-orthogonality properly.
@@
Figure 2: Spin profile of single site excitations $\langle S_{z}(x)\rangle$ versus site $x$. (a) Excitation profiles for the lowest energy single-site state, where a single tensor is different from the ground state. The numbers beside each curve indicate the cutoff $\lambda^{2}$ for the pretruncation step; these cutoffs result in $\chi_{\text{tr}}=8,24,46$, and $90$ for the curves top to bottom. (b) Comparison of excitation profile for states with and without restriction to the left null space, both with $\lambda^{2}=10^{-10}$.
@@
Because the ground state is gapped, connected correlations decay as $e^{-|j-j^{\prime}|/\xi}$ and both kernels are therefore exponentially banded. In practice we take $L_{c}\leq j-j^{\prime}\leq L_{c}$ with $L_{c}\simeq 150$; beyond that threshold every entry is below about $10^{-12}$ and can be safely discarded. The calculation of the $H$ array of size $N_{\alpha}^{2}L_{c}$ is the dominant effort in SBEA. If the Hamiltonian MPO is in infinite uniform $\ldots HHH\ldots$ form, then the contractions are particularly simple, since the edge tensors don’t depend on the site. In that case for each $B_{\alpha}$ we start from the left, and progressive contract towards the right. For each site, there is a final contraction of all the $B_{\alpha^{\prime}}$. Then the calculation time is $O[(N_{\alpha}^{2}+N_{\alpha}K)L_{c}\chi^{3}]$, where $K$ is the MPO bond dimension. The calculation time for the different curves in Fig. 3 varies primarily as $N_{\alpha}^{2}$. This calculation time is identical to that of a finite-system DMRG calculation on a system of length $L_{c}$, except the parenthetical factor would be replaced by an effective number of sweeps times the number of Lanczos steps at each step, i.e. still similar to a single DMRG ground state calculation.